# Landauer Principle and Thermodynamics of Computation

### v1

@@ Abstract
According to the Landauer principle, any logically irreversible process accompanies entropy production, which results in heat dissipation in the environment. Erasing of information, one of the primary logically irreversible processes, has a lower bound on heat dissipated into the environment, called the Landauer bound (LB). However, the practical erasure processes dissipate much more heat than the LB. Recently, there have been a few experimental investigations to reach this bound both in the classical and quantum domains. There has also been a spate of activities to enquire about this LB in finite time, with finite-size heat baths, non-Markovian and nonequilibrium environment in the quantum regime where the effects of fluctuations and correlation of the systems with the bath can no longer be ignored. This article provides a comprehensive review of the recent progress on the Landauer bound, which serves as a fundamental principle in the thermodynamics of computation. We also provide a perspective for future endeavors in these directions.
Furthermore, we review the recent exploration toward establishing energetic bounds of a computational process. We also review the thermodynamic aspects of error correction, which is an indispensable part of information processing and computations. In doing so, we briefly discuss the basics of these fields to provide a complete picture.
@@
$\bullet$ LP $\rightarrow$ Landauer Principle
@@ I Introduction
John von Neumann, in his 1948 lectures, posed a fundamental question regarding whether a computer operating at temperature $T$ must necessarily dissipate heat. The intricate details of this concept were later formalized in his book Theory of Self-Reproducing Automata, which was completed in 1966 by Arthur W. Burks Von Neumann et al. (1966). Building upon this idea, Landauer famously asserted that “real-world computation involves thermodynamic costs” and emphasized the substantial implications of this principle Landauer (1961, 1991). Further support for this notion came from Brillouin’s thought experiment Brillouin (1962), which validated the connection between computation and thermodynamics, albeit with some error probability. The man-made computers and even all naturally occurring processes like biological computers have thermodynamic costs. It is quite fascinating to analyze the difference in the thermodynamic cost of the naturally occurring process and the artificial ones that are created. Translation of ribonucleic acid (RNA) into amino acids is one such natural biological process where one encounters energy costs for the execution of the process. These biological processes are thermodynamically more efficient than the artificial ones Kempes et al. (2017).
@@
These examples of artificial and natural systems hinge on the deep connection between computation and thermodynamics. The connection between thermodynamics and a logically irreversible process gets most prominent in the context of Maxwell’s demon (MD) Maxwell (1871); Maruyama et al. (2009). Landauer (later Bennett) argued that one must pay an entropic cost that gets dumped in the environment as heat while performing a logically irreversible process that erases or throws away information. This argument played the pivotal role in exorcising the MD and thereby saving the second law of thermodynamics. It also established that information is physical. There have been several explorations to model MD in physical systems and study the Landauer principle (LP). However, they mostly belong to equilibrium statistical physics. But when we approach to the miniaturized domain, the systems are, in general, highly non-equilibrium in their form. The major breakthroughs in non-equilibrium statistical physics Zwanzig (2001); Kubo et al. (2012); Jarzynski (1997a, b) allow us to analyze the thermodynamic behavior of systems that are arbitrarily far from their equilibrium, and even for systems that undergo arbitrary external driving. With the advent of non-equilibrium statistical physics, the researchers have shown a keen interest to analyze the thermodynamic cost for the erasure process in this domain which includes finite time, finite size, open quantum systems, and so on Goold et al. (2015); Khoudiri et al. (2025); Lorenzo et al. (2015); Proesmans et al. (2020a); Miller et al. (2020); Esposito and Van den Broeck (2011); Mandal and Jarzynski (2012); Helms and Limmer (2022); Ray and Crutchfield (2023); Kuang et al. (2022); Proesmans and Bechhoefer (2021); Barnett and Vaccaro (2013); Berta et al. (2018); Henao and Uzdin (2023); Barato and Seifert (2013); Mandal et al. (2013); Deffner and Jarzynski (2013); Barato and Seifert (2014); Strasberg et al. (2014).
@@
In nature, copying information is a fundamental mechanism within natural systems Watson and Crick (1953); Hopfield (1974); Ouldridge et al. (2017); Horowitz and England (2017a). Yet, replication is inherently prone to error. The accuracy of a copy thus relies on its accurate reproduction. It can be quantified by counting the wrongly copied bits while execution. Though the error can be reduced at the macro level, at the molecular level, perfect copying is not achievable due to thermal fluctuations. It is the primary source of error at the molecular level. The replication process is limited by the thermal noise, so it must be interpreted in terms of thermodynamics as proposed by von Neumann Von Neumann (1956). This raises an important question: whether one can develop a connection between the errors that occur when copying remains.
@@
Generally, while executing a copying process, it has to undergo various intermediate steps to control the accuracy and speed of the process. This is true for artificial as well as natural scenarios Johnson (1993). If we try to explain the errors from thermodynamic laws (second law), one must account for the fact that the copying process is repeated cyclically rather than occurring as a single-shot operation Bennett (1982). The works Sartori and Pigolotti (2015); Korepin and Terilla (2002) forge a profound link between thermodynamics and informational errors, demonstrating that errors arising in copying protocols are intrinsically connected to thermodynamic observables, which characterize the errors.Another aspect to connect thermodynamics and information theory is built on the basis of using entanglement in quantum systems Horodecki et al. (2001); Popescu and Rohrlich (1997); Rohrlich (2001).
@@
In this article, we review the recent developments in Landauer erasing in various contexts, including the most recent experimental demonstrations, thermodynamic aspects, and cost analysis of computation and error correction.
@@
Aiming to highlight recent advancements and fundamental insights into Landauer’s principle (LP) and the thermodynamics of computation, this review addresses several key aspects spanning both theoretical and experimental fronts. We begin in Sec. II with a discussion of the generalized LP. Sec. III reviews significant recent developments concerning LP, including scenarios involving finite time, finite-size systems, and non-Markovian reservoirs. In Sec. IV, we explore the role of Landauer’s limit (LL) in computational processes, followed by Sec. V, which presents an overview of recent experimental efforts aimed at realizing and validating LL in practical systems.
Consequently, we move towards exploring the thermodynamic ramifications and implications for computation and error correction. Sec. VI delves into reversible computational models, laying the groundwork for a broader discussion in Sec. VII on the thermodynamics of computational paradigms, including finite state automata and Turing machines. In Sec. VIII, we critically examine the thermodynamic consistency of error-correcting codes. Prior to concluding and outlining potential future research directions in Sec. X, we briefly address a range of complementary topics under the umbrella of miscellaneous discussions in Sec. IX, such as viewing the computer as a heat engine, the thermodynamics of algorithms, and Landauer bound in switching protocols.
@@
Though the review covers an extensive amount of works in the context of LP and thermodynamics of computation, there are some related topics and works, which are not discussed, that are worth mentioning. For exmaple, while most of the aspects of the LP have been reviewed here, some of the left-out aspects in this domain are like the role of LP in gravity Bormashenko (2019a); Haranas et al. (2021); Daffertshofer and Plastino (2007), relativity Herrera (2020), quantum field theory Xu et al. (2022), many-body phenomena Bonança (2023); Parrondo (2001), material Zivieri (2022).
In this review, we have considered mainly two fundamental computational models: Turing machine and Finite state machine, namely, for detailed thermodynamic analysis. Some aspects, like LB in analog computers and algorithmic thermodynamics, have not been discussed here. Interested readers can go through Diamantini et al. (2016) for the LB analysis in the analog system and Baez and Stay (2012) for algorithmic thermodynamics, which in turn allows one to apply the laws and techniques of thermodynamics for studying algorithmic information theory; in contrast, we have focused on the thermodynamic cost of algorithms in Sec. IX.
In what follows, we briefly deal with computational complexity when we estimate the energetic cost of computation. However, a somewhat related direction, the fundamental limitation on the computability of the physical process, is not discussed in this review. Interested readers can have a look at the following seminal articles Pour-El and Richards (1982); Moore (1990); Lloyd (2000, 2017). Similarly, the thermodynamics of controlled systems is not covered. Please go through Touchette and Lloyd (2004, 2000); Barato and Seifert (2017); Sagawa and Ueda (2008, 2012); Wilming et al. (2016); Large and Large (2021); Gingrich et al. (2016); Horowitz and England (2017b) for details.
@@ II Landauer’s Principle
The search for the development of more impactful computing circuits leads to the question: What would be the physical limitation of this process? Rolf Landauer, in the year 1961, in his seminal work Landauer (1961) proposed an important limit to the conjecture surfaced by von Neumann Von Neumann et al. (1966), which is coined as “Landauer bound” (LB). This is a physical principle that provides the lower bound on the consumption of a computation process, i.e., an irreversible process in the computer dissipates a minimum amount of heat per bit Bennett (2003), which is expressed as
$\Delta Q\geq k_{B}T\ln 2,$ (1)
@@
where $k_{B}$ is the Boltzman constant and $T$ is the temperature at which the computation process operates. The bound at room temperature is approximately 0.018 $eV$ (2.9 x $10^{-21}$ J), whereas the modern world computers use a million times more energy per operation Lambson et al. (2011a); Moore (2012).
@@
Landauer Principle (LP) though widely accepted has faced various challenges and objections Earman and Norton (1998, 1999); Shenker (1998); Maroney (2005); Norton (2005, 2011); Bennett (2003). The prime objections to LP were put forth by Earman and Norton Earman and Norton (1999), where they have argued that LP is dependent on the second law of thermodynamics, it can be considered to be either unnecessary or insufficient for “exorcism of Maxwell’s Demon”. The other objections over LP are generally of three kinds Bennett (2003): (a) The principle fails, as thermodynamic quantities such as heat and work are fundamentally unrelated to mathematical constructs like logical reversibility. As a result, drawing direct parallels between them lacks any meaningful basis (as it was thought of to be a memory erasure process at that time). (b) In all cases of data-processing operations, there is a dissipation of at least $k_{B}T\ln 2$ amount of energy irrespective of the condition whether it is logically reversible or not. (c) Another reason for objection is that, in principle, logically irreversible operations can be engineered in a thermodynamically reversible way.
@@
Attempts have been made to demonstrate LP with precision. Piechocinska Piechocinska (2000) offered proof of LP with the help of statistical mechanics. The proof had a particular physical model assumption. The notable advancement was provided in the work Turgut (2009); Ladyman et al. (2007, 2008); Cao and Feito (2009); Leff and Rex (2002), where it provides a generalized version of the LP without a particular physical model assumption. The generalization of the LP surfaced in Vaccaro and Barnett (2011), where it has been shown that the information erasure will cause an increase in the entropy of the environment with no energy cost but the cost can be attributed to angular momentum of a spin-reservoir. In fact, angular momentum is exploited rather than energy from a spin-reservoir to erase the memory and it is consistent with the second law of thermodynamics.
@@
In the low-temperature limit, quantum effects play an important role in the analysis of the thermodynamic aspect of the erasure principle. The validity of the erasure principle in this limit has been challenged Allahverdyan and Nieuwenhuizen (2000); Nieuwenhuizen and Allahverdyan (2002); Hörhammer and Büttner (2005); Hilt and Lutz (2009). It has been claimed that due to the presence of entanglement between the system and the environment LP is broken, which implies information is erased while heat is absorbed Allahverdyan and Nieuwenhuizen (2001); Hörhammer and Büttner (2008); Cápek and Sheehan (2005); Maruyama et al. (2009) which counters the LP. Lutz et al. Hilt et al. (2011), tackled this quantum conundrum and demonstrated that LP remains valid regardless of the specific nature of the interaction between the system and its environment. Their findings provided significant evidence supporting the applicability of LP in quantum settings.
@@
To address the need for a theory with minimal assumptions in both regimes, approaches are considered that derive the LP without relying explicitly on the second law of thermodynamics Shizume (1995); Piechocinska (2000); Sagawa and Ueda (2009, 2011). Reeb’s work Reeb and Wolf (2014) is on this line of thought. The assumptions that are considered for the analysis are minimal based on the benchmark works. The assumptions are: (a) In the process, the system and the reservoir belong to a Hilbert space, (b) the reservoir is initially in a thermal state, (c) the system and the reservoir will be initially uncorrelated, and (d) the evolution process is unitary.
@@
With this minimum condition, a sharpened equality version of LP has been derived for a system $\mathcal{S}$ and a reservoir $R$ as
$\beta\Delta Q=\Delta S+I(\mathcal{S}^{\prime}:R^{\prime})+S(\rho_{R}^{\prime}||\rho_{R})\geq\Delta S.$ (2)
@@
Here $\prime$ denotes the final states, $\Delta S$ is the change of von Neumann entropy of the system, $I(\mathcal{S}^{\prime}:R^{\prime})$ describes the mutual information and quantifies the correlation of the system with the bath, and $\rho_{R}$ denotes the reservoir state. $\beta=1/k_{B}T$ denotes the inverse temperature of the bath. $S(\rho_{R}^{\prime}||\rho_{R})$ quantifies the free energy increase in the bath, where $S(\rho||\rho^{\prime})=\text{Tr}\,\rho\,\ln\,\rho-\text{Tr}\,\rho\,\ln\,\rho^{\prime}$ is the quantum relative entropy between $\rho,\,\rho^{\prime}$.
@@
The setup mentioned above can be generalized by considering an initial correlation in the process. So the assumption gets modified as: (a) In the process, the system, reservoir, and memory are in a joint quantum state initially, (b) the reservoir is initially in a thermal state, (c) the heat exchange and the entropy of the system are expounded on the marginal states, and (d) the evolution is processed by an unital positive trace-preserving map. In this generalized setup, where the system, reservoir, and memory are initially correlated, the standard Landauer bound can be modified significantly. Initial quantum or classical correlations can effectively reduce the minimal heat dissipation required for information erasure. Since entropy and heat are evaluated from marginal states, part of the entropy change may be absorbed by shared correlations rather than dissipated as heat. When the evolution is governed by a unital, trace-preserving map, which does not reduce entropy on its own, the role of initial correlations becomes even more prominent.
@@
In the present-day protocols where the LP plays a vital role, Reeb’s version of the erasure process is considered for the analysis. Even with the minimal assumption of Reeb, it is clear that one requires dissipative dynamics to explore the LP. However, it is a difficult task in pure Hamiltonian dynamics where the evolution is for the system only. It has been conveyed in Holtzman et al. (2021) that it is possible to implement an erasure bit with no thermodynamic cost using Hamiltonian dynamics if one has the information of the system with infinite accuracy. In this case, it corresponds to the energy of the system being known with infinite accuracy.
@@
So far, we have discussed LB for a non-zero finite temperature $T$ of the environment. What if the temperature of the environment tends to zero? In this limit (i.e., $T\rightarrow 0$) the bound in Eq. (1) is trivial. It states that $Q\geq 0$ when $T=0$ as the bath is in the ground state. A non-trivial bound for this condition has been proposed in Timpanaro et al. (2020). There, the authors have assumed that initially the environment is in a thermal state, as in the original case of LEP. The state of the system or the type of system-environment interaction is considered to be general. The improved bound is always greater than Eq. (1) and coincides with it when $T$ is high. It is derived based on two principles, the positivity of mutual information and the maximum entropy principle Guiasu and Shenitzer (1985); Wu (2012); Pressé et al. (2013), namely, and can be expressed in terms of the equilibrium heat capacity of the bath. It is interesting to explore what the modifications will be on the bound when the system and environment are initially correlated for the limit $T\rightarrow 0$.

### v2

# thanks: Contributed equallythanks: Contributed equally

@@ Abstract
According to the Landauer principle, any logically irreversible process accompanies entropy production, which results in heat dissipation in the environment. Erasing of information, one of the primary logically irreversible processes, has a lower bound on heat dissipated into the environment, called the Landauer bound (LB). However, the practical erasure processes dissipate much more heat than the LB. Recently, there have been a few experimental investigations to reach this bound both in the classical and quantum domains. There has also been a spate of activities to enquire about this LB in finite time, with finite-size heat baths, non-Markovian and nonequilibrium environments in the quantum regime where the effects of fluctuations and correlation of the systems with the bath can no longer be ignored. This article provides a comprehensive review of the recent progress on the Landauer bound, which serves as a fundamental principle in the thermodynamics of computation. We also provide a perspective for future endeavors in these directions.
Furthermore, we review the recent explorations toward establishing energetic bounds of a computational process. We also discuss the thermodynamic aspects of error correction, which is an indispensable part of information processing and computations. In doing so, we briefly discuss the basics of these fields to provide a complete picture.
@@
$\bullet$ LP $\rightarrow$ Landauer’s Principle
@@ I Introduction
John von Neumann, in his 1948 lectures, posed a fundamental question on whether a computer operating at temperature $T$ must necessarily dissipate heat. The details of this lectures were later formalized in his book Theory of Self-Reproducing Automata, completed in 1966 by Arthur W. Burks Von Neumann et al. (1966). In 1961, Landauer asserted that “real-world computation involves thermodynamic costs” and emphasized the implications of the same Landauer (1961, 1991). The man-made computers and even all naturally occurring processes like biological computers have thermodynamic costs. It is quite fascinating to analyze the difference in the thermodynamic cost of the naturally occurring process and the artificial ones that are created. Translation of ribonucleic acid (RNA) into amino acids is one such natural biological process where one encounters energy costs for the execution of the process. These biological processes are thermodynamically more efficient than the artificial ones Kempes et al. (2017).
@@
These examples of artificial and natural systems hinge on the deep connection between computation and thermodynamics. The connection between thermodynamics and a logically irreversible process gets most prominent in the context of Maxwell’s demon (MD) Maxwell (1871); Maruyama et al. (2009); Junior et al. (2025). Landauer (later Bennett) argued that one must pay an entropic cost that gets dumped into the environment as heat while performing a logically irreversible process that erases or throws away information. This argument played the pivotal role in exorcising the MD and thereby saving the second law of thermodynamics. It also established that information is physical. There have been several explorations to model MD in physical systems and study Landauer’s principle (LP). However, they mostly belong to equilibrium statistical physics. But when we approach to the miniaturized domain, the systems are, in general, out of equilibrium. The major breakthroughs in non-equilibrium statistical physics Zwanzig (2001); Kubo et al. (2012); Jarzynski (1997a, b) allow us to analyze the thermodynamic behavior of systems that are arbitrarily far from their equilibrium, and even for systems that undergo arbitrary external driving. With the advent of non-equilibrium statistical physics, the researchers have shown a keen interest to analyze the thermodynamic cost for the erasure process in this domain which includes finite time, finite size, non-Markovian (NM) open quantum dynamical effects, and so on Goold et al. (2015); Khoudiri et al. (2025); Lorenzo et al. (2015); Proesmans et al. (2020a); Miller et al. (2020); Esposito and Van den Broeck (2011); Mandal and Jarzynski (2012); Helms and Limmer (2022); Ray and Crutchfield (2023); Kuang et al. (2022); Proesmans and Bechhoefer (2021); Barnett and Vaccaro (2013); Berta et al. (2018); Henao and Uzdin (2023); Barato and Seifert (2013); Mandal et al. (2013); Deffner and Jarzynski (2013); Barato and Seifert (2014); Strasberg et al. (2014); Ma et al. (2022); Mondal et al. (2023).
@@
In nature, copying information is a fundamental mechanism within natural systems Watson and Crick (1953); Hopfield (1974); Ouldridge et al. (2017); Horowitz and England (2017a). Yet, replication is inherently prone to error. The accuracy of a copy thus relies on its accurate reproduction. It can be quantified by counting the wrongly copied bits while execution. Though the error can be reduced at the macro level, but at the molecular level, perfect copying is not achievable due to thermal fluctuations. It is the primary source of error at the molecular level. The replication process is limited by the thermal noise, so it must be interpreted in terms of thermodynamics as proposed by von Neumann Von Neumann (1956). This raises an important question: Can one establish a fundamental connection between the thermodynamic cost and the errors incurred during the information copying process?
@@
Generally, while executing a copying process, it has to undergo various intermediate steps to control the accuracy and speed of the process. This is true for artificial as well as natural scenarios Johnson (1993). If we try to explain the errors from thermodynamic laws (the second law), one must account for the fact that the copying process is repeated cyclically rather than occurring as a single-shot operation Bennett (1982). The works Sartori and Pigolotti (2015); Korepin and Terilla (2002) forge a profound link between thermodynamics and informational errors, demonstrating that errors arising in copying protocols are intrinsically connected to thermodynamic observables, which characterize the errors.
Understanding the thermodynamic cost of computation is essential in both classical and quantum computation, as it reveals how the physical laws fundamentally constrain computational efficiency. Core models such as reversible computation, Turing machines, and finite automata offer deep theoretical insights into the thermodynamic cost of computational models. Reversible computing frameworks, in principle, allow for vanishingly low energy dissipation by preserving information throughout the process, yet practical limitations inevitably lead to non-negligible thermodynamic costs. Similarly, reversible Turing machines can asymptotically approach this ideal limit, though operations such as memory resetting and tape erasure introduce unavoidable energy losses. Finite automata, despite their relative simplicity, capture the essence of how logical state transitions and memory utilization contribute to entropy production. Collectively, these models provide a rigorous foundation for exploring the intrinsic trade-offs among energy consumption, computational speed, and information preservation in physically realizable computing systems.
@@
In this article, we review the recent developments on LP in various contexts, including the most recent experimental demonstrations, thermodynamic aspects, and cost analysis of computational models and error correction.
@@
Aiming to highlight recent advancements and fundamental insights into LP and the thermodynamics of computation, this review addresses several key aspects spanning both theoretical and experimental fronts. We begin in Sec. II with a discussion of the generalized LP. Sec. III reviews significant recent developments on LP, including scenarios involving finite time dynamics, finite-size and non-Markovian reservoirs. In Sec. IV, we explore the role of Landauer limit (LL) in computational processes, followed by Sec. V, which presents an overview of recent experimental efforts aimed at realizing and validating LL in practical systems.
Consequently, we move towards exploring the thermodynamic ramifications and implications for computation and error correction. Sec. VI delves into reversible computational models, laying the groundwork for a broader discussion in Sec. VII on the thermodynamics of computational paradigms, including finite state automata and Turing machines. In Sec. VIII, we discuss the thermodynamic consistency of error-correcting codes. Prior to concluding and outlining potential future research directions in Sec. X, we briefly address a range of complementary topics such as viewing the computer as a heat engine, the thermodynamics of algorithms, and the Landauer bound (LB) in switching protocols, under the umbrella of miscellaneous discussions in Sec. IX.
@@
Though the review covers an extensive amount of works in the context of LP and thermodynamics of computation, there are some related topics and work, that are not discussed here, but are worth mentioning. For example, while most of the aspects of the LP have been reviewed here, some of the left-out aspects in this domain are like the role of LP in gravity Bormashenko (2019a); Haranas et al. (2021); Daffertshofer and Plastino (2007), relativity Herrera (2020), quantum field theory Xu et al. (2022), many-body phenomena Bonança (2023); Parrondo (2001), material Zivieri (2022).
In this review, we have considered mainly two fundamental computational models: the Turing machine and the Finite state machine, namely, for detailed thermodynamic analysis. Some aspects, like LB in analog computers and algorithmic thermodynamics, have not been discussed here. Interested readers can go through Diamantini et al. (2016) for the LB analysis in the analog system and Baez and Stay (2012) for algorithmic thermodynamics, which in turn allows one to apply the laws and techniques of thermodynamics for studying algorithmic information theory; however, we have touched upon the thermodynamic cost of algorithms briefly in Sec. IX.
In what follows, we briefly deal with computational complexity when we estimate the energetic cost of computation. However, a somewhat related direction, the fundamental limitation on the computability of the physical process, is not discussed in this review. Interested readers can have a look at the following seminal articles Pour-El and Richards (1982); Moore (1990); Lloyd (2000, 2017). Similarly, the thermodynamics of controlled systems is not covered. Interested readers can go through Touchette and Lloyd (2004, 2000); Barato and Seifert (2017); Sagawa and Ueda (2008, 2012); Wilming et al. (2016); Large and Large (2021); Gingrich et al. (2016); Horowitz and England (2017b) for details.
@@ II Landauer’s Principle
The search for the development of more impactful computing circuits leads to the question: What would be the physical limitation of this process? Rolf Landauer, in the year 1961, in his seminal work Landauer (1961) proposed an important limit to the posed question by von Neumann Von Neumann et al. (1966). Landauer stated that an irreversible process, where information is either deleted or thrown away, dissipates a minimum amount of energy per bit Bennett (2003), which is dumped as heat into the environment, increasing its entropy. This minimum amount of heat dissipated to erase one bit of information, known as “Landauer bound” (LB), is given by
$Q\geq k_{B}T\ln 2,$ (1)
@@
where $k_{B}$ is the Boltzman constant and $Q$ is the heat dumped into the environment at temperature $T$. The bound at room temperature is approximately 0.018 $eV$ (2.9 x $10^{-21}$ J), whereas the modern world computers use a billion times more energy Lambson et al. (2011a); Moore (2012).
In 1867, Maxwell proposed a thought experiment, which he later elaborated on in his book “Theory of heat” Maxwell (1871), where a demon has a chamber of ideal gas in equilibrium at some fixed temperature. The demon creates a partition between the chamber and allows the molecules to cross the partition from, say, left to right if their velocity is greater than some threshold value. The molecules in the right can only go to the left if their velocities are less than the threshold value. Thus, the demon can create a temperature difference, which can further be exploited to extract work in principle, but without performing any work. This contradicts the second law of thermodynamics, as one can extract work from a single temperature source. Szilard proposed a simplified version of this thought experiment with a single gas molecule known as Szilard’s engine in 1929. This thought experiment puzzled the scientific communities for almost a century till the discovery of LP in 1961 Landauer (1961). The exorcism of MD can be done by virtue of the fact that the demon needs at least the amount of energy it has extracted earlier, on average, to erase their memory, before it can work in the next cycle. Brillouin’s argument to reconcile MD with the second law was that the demon must pay an entropic cost to acquire information, which he dubbed as negentropy Brillouin (1962). It was Bennet who showed in 1982 Bennett (1982) that the information gain and storage can be done reversibly in principle, and confirmed that only the erasure step must bear the entropic cost. Thus, the LP successfully exorcises the MD.
@@
Though widely accepted for reconciling MD with the second law Ladyman et al. (2007); Bennett (2003), Landauer’s Principle (LP) has faced various criticisms and objections Earman and Norton (1998, 1999); Shenker (1998); Maroney (2005); Norton (2005, 2011); Bennett (2003). The prime objections to LP were put forth by Earman and Norton Earman and Norton (1999). They argued that LP is dependent on the second law of thermodynamics; it can be considered to be either unnecessary or insufficient for “exorcism of Maxwell’s Demon”. In Sagawa and Ueda (2008); Cao and Feito (2009) it has also been argued recently that LP is a consequence of the second law. The other objections over LP are generally of three kinds Bennett (2003): (a) According to many, at that time at least, thermodynamic quantities such as heat and work are fundamentally unrelated to mathematical concepts like logical irreversibility. Therefore, drawing direct parallels between them seems unfounded. (b) In all cases of data-processing operations, there is a dissipation of at least $k_{B}T\ln 2$ amount of energy irrespective of the condition whether it is logically reversible or not. In a realistic physical scenario, it is even much more. (c) Another reason for objection is that, in principle, even logically irreversible operations can be engineered in a thermodynamically reversible way. In Bennett (2003), the author discussed in detail and refuted all these criticisms.
@@
Attempts have been made to establish LP rigorously. Piechocinska Piechocinska (2000) offered a proof of LP with the help of statistical mechanics. The proof assumes a particular physical model. The notable advancements have been done in Turgut (2009); Ladyman et al. (2007, 2008); Cao and Feito (2009); Leff and Rex (2002), where LP has been generalized without assuming a particular physical model. A notable generalization of the LP surfaced in Vaccaro and Barnett (2011), where it has been shown that the information erasure will cause an increase in the entropy of the environment with no energy cost, but the cost can be attributed to the angular momentum of a spin-reservoir. In fact, angular momentum is exploited rather than energy from a spin-reservoir to erase the memory, and it is consistent with the second law of thermodynamics.
@@
With rapid technological advancements and our ever increasing grasp on the miniaturized scale, probing and controlling the small scale systems that exhibits quantum traits are becoming more and more possible. In this regime, where we can no longer ignore the fluctuations, quantum correlations between the system and the environment, the quantum effects bring fundamental challenges in the analysis of memory erasure. The validity of the erasure principle in this limit has been questioned Allahverdyan and Nieuwenhuizen (2000); Nieuwenhuizen and Allahverdyan (2002); Hörhammer and Büttner (2005); Hilt and Lutz (2009). It has been claimed that due to the presence of entanglement between the system and the environment LP is violated, which implies information is erased while heat is absorbed Allahverdyan and Nieuwenhuizen (2001); Hörhammer and Büttner (2008); Cápek and Sheehan (2005); Maruyama et al. (2009). Hilt et al. Hilt et al. (2011), have tackled this quantum conundrum and demonstrated that LP remains valid regardless of the specific nature of the interaction between the system and its environment. Their findings have provided significant evidence supporting the applicability of LP in the quantum settings.
@@
To address the need for a theory with minimal assumptions, approaches have been considered that derive the LP without relying explicitly on the second law of thermodynamics in both classical and quantum regimes Shizume (1995); Piechocinska (2000); Sagawa and Ueda (2009, 2011). Reeb and Wolf’s work Reeb and Wolf (2014) is the most significant one along this direction in quantum setups. They have considered the following minimal set of assumptions: (a) In the process, the system and the reservoir belong to corresponding Hilbert spaces. (b) The reservoir is initially in a thermal state. (c) The system and the reservoir are initially uncorrelated. (d) The evolution process is unitary on the joint system-resrvoir state.
@@
With this minimal setup, a sharpened equality version of LP has been derived for a system $\mathcal{S}$ and a reservoir $R$, for the dissipated heat $Q$ to the environment as
$\beta Q=\Delta S+I(\mathcal{S}^{\prime}:R^{\prime})+S(\rho_{R}^{\prime}||\rho_{R})\geq\Delta S.$ (2)
@@
Here $\prime$ denotes the final states, $\Delta S$ is the decrease in von Neumann entropy of the system, $I(\mathcal{S}^{\prime}:R^{\prime})$ describes the mutual information which quantifies the correlation of the system with the bath Nielsen and Chuang (2002), and $\rho_{R}$ denotes the reservoir state. $\beta=1/k_{B}T$ denotes the inverse temperature of the bath. $S(\rho_{R}^{\prime}||\rho_{R})$ quantifies the free energy increase in the bath, where $S(\rho||\rho^{\prime})=\text{Tr}\,[\rho\,\ln\,\rho]-\text{Tr}\,[\rho\,\ln\,\rho^{\prime}]$ is the quantum relative entropy between $\rho,\,\rho^{\prime}$ Nielsen and Chuang (2002).
@@
They further generalized the setup Reeb and Wolf (2014) by considering an initial correlation in the process, which modifies the assumption as follows: (a) In the process, the system, reservoir, and memory are in a joint quantum state initially, (b) the marginal state of the reservoir is initially thermal, (c) the heat and entropy exchange are expounded on the marginal states, and (d) the evolution is processed by an unital positive trace-preserving map. In this generalized setup, where the system, reservoir, and memory are initially correlated, the standard Landauer bound can be modified significantly. Initial quantum or classical correlations can effectively reduce the minimal heat dissipation required for information erasure. Since entropy and heat are evaluated from marginal states, part of the entropy change may be absorbed by shared correlations rather than dissipated as heat. When the evolution is governed by a unital, trace-preserving map, which does not reduce entropy on its own, the role of initial correlations becomes even more prominent.
@@
Even with the minimal assumption of Reeb and Wolf, it is considered that the system is governed by non-unitary dynamics. However, it is a difficult task to explore erasing in pure Hamiltonian dynamics for the system. It has been conveyed in Holtzman et al. (2021) that it is possible to implement a bit-erasure with no thermodynamic cost using Hamiltonian dynamics if one has the information of the system with infinite accuracy. In this case, it corresponds to the energy of the system being known with infinite accuracy. This requires diverging computational resources as we will discuss later in Sec. III B.
@@
So far, we have discussed LB for a non-zero finite temperature $T$ of the environment. What if the temperature of the environment tends to zero? In this limit (i.e., $T\rightarrow 0$) the bound in Eq. (1) is trivial. It states that $Q\geq 0$ when $T=0$ as the bath is in the ground state. A non-trivial bound for this condition has been proposed in Timpanaro et al. (2020). There, the authors have assumed that initially the environment is in a thermal state, as in the original case of LP. The state of the system or the type of system-environment interaction is considered to be general. The improved bound is always greater than Eq. (1) and coincides with it when $T$ is high. It is derived based on two principles, the positivity of mutual information and the maximum entropy principle Guiasu and Shenitzer (1985); Wu (2012); Pressé et al. (2013), namely, and can be expressed in terms of the equilibrium heat capacity of the bath. It is interesting to explore what the modifications will be on the bound when the system and environment are initially correlated for the limit $T\rightarrow 0$.